OpenStudy (abmon98):

f(x) 7cosx-24sinx Given that f(x)=Rcos(x+a) where R>=0, 0<=a<=1/2pi and x and a are measured in radians. Find R and show that a=1.29 to 2 d.p Hence Write down the minimum value of f(x) The value of x in the interval 0<=x<=2pi Find the smallest possible two positive values of x for which 7cosx-24sinx=10

ganeshie8 (ganeshie8):

where are you stuck

ganeshie8 (ganeshie8):

you have been using `acosx+bsinx` formula right ?

OpenStudy (abmon98):

yes i do use this formula @ganeshie8 7cosx-24sinx=Rcosxcosa+RSinxsina 7=Rcosa -24=Rsina Tan=-24/7 R=sqrt(7)^2+(-24)^2=25

ganeshie8 (ganeshie8):

Yes ! \[7\cos x-24\sin x = 25\cos\left(x - \tan^{-1}(\frac{-24}{7})\right)\]

OpenStudy (abmon98):

So how is a=1.29

ganeshie8 (ganeshie8):

whats the value of \(\tan^{-1}(\frac{-24}{7})\) ?

OpenStudy (abmon98):

oh i get how to do it now tan-1(-24/7)=-73.73979529*pi/180=-1.287 rounded to 2d.p -1.29

ganeshie8 (ganeshie8):

Yep !

ganeshie8 (ganeshie8):

\[7\cos x-24\sin x = 25\cos\left(x - \tan^{-1}(\frac{-24}{7})\right)\] \[= 25\cos\left(x - (-1.29)\right)\] \[= 25\cos\left(x +1.29\right)\]

OpenStudy (abmon98):

Thank you @ganeshie8 but how can i solve to find the value of x in the interval 0<=x<=2pi to get this minimum value

OpenStudy (dls):

\[\LARGE -25 \leq 7cosx x−24\sin x \leq 25\] clearly minimum value is -25